How do I add these rational expressions? 2 / y - 1 + 2 / y + 1 ?
Answer by fractalier(2101)

2 / (y - 1) + 2 / (y + 1)
For this, you first determine the lowest common denominator and then multiply each fraction in such a way that the denominators becomes that LCD...in this case, the LCD is (y - 1)(y + 1)...
We multiply the first by (y + 1) / (y + 1)
and the second by (y - 1) / (y - 1) and get
2(y + 1) / (y - 1)(y + 1) + 2(y - 1) / (y - 1)(y + 1) =
[2(y + 1) + 2(y - 1)] / (y - 1)(y + 1) =
4y / (y - 1)(y + 1)
The easiest way is to apply the criss-cross method...
A/B + C/D = (AD + BC) / BD